OpenStudy (anonymous):

2x - 6y = 5 x + y = 2 Solve the system of equations. i got (17, -15)??

OpenStudy (anonymous):

\(2x - 6y = 5\) + \(-2(x + y) = -2(2)\) ============== \(0x -8y = 1\) I don't think what you have there is correct.

OpenStudy (anonymous):

is it (17/8, -1/8)?

OpenStudy (anonymous):

That's correct.